OpenStudy (anonymous):

The total area of Egypt is 995,450 mi2. 19/20 of the area is desert. The remaining land is used for growing crops and is where civilizations are located. What is the area of the land that is desert? What is the area of the land that is usable? Write all of the given information that will help you solve this problem. Explain, using complete sentences, how you will solve this problem.Write the mathematical expression or expressions that you will use to solve this problem. Carry out the operations carefully. Show all of your work. Check your results.

OpenStudy (tkhunny):

You have a list. Get started! Explain, using complete sentences, how you will solve this problem. Write the mathematical expression or expressions that you will use to solve this problem. Carry out the operations carefully. Show all of your work. Check your results.

OpenStudy (anonymous):

I dont know how to though. Thats what I need answered. @tkhunny

OpenStudy (tkhunny):

You must try something. Give the first one a go!

OpenStudy (anonymous):

I dont know how to solve it .... @tkhunny I've tried a million times and it gets me nowhere.

OpenStudy (tkhunny):

You have the total area of Egypt. You have a ratio to split the land into two parts. How shall that be accomplished?

OpenStudy (anonymous):

I dont know @tkhunny

OpenStudy (tkhunny):

That's just not good enough. You absolutely must do better than that. It won't fall out of the sky. You must give it some thought. What section are you studying? Ratios? Percentages? Sets?

OpenStudy (anonymous):

Im no longer on this. It has to be turned in by January 1st. Im now on planes.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Fine. What section WERE you studying when you were given this problem? You must help, here. I must not do it for you.

OpenStudy (anonymous):

I was studying Fractions at this time. @tkhunny Im homeschooled so i dont go by sections like public schools.

OpenStudy (tkhunny):

It matters not at all what path you take. What is important is that you understand what you are doing. If we cut a pie into 10 pieces, and I eat 1/5 of the pie, how many pieces did I eat?

OpenStudy (anonymous):

I dont understand how you get 1/5 out of 10 pieces.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Very good. This is useful information. Ever hear of "Equivalent Fractions"? Do you know that these fractions are the SAME? 1/3 = 2/6 = 15/45 = 70/210

OpenStudy (anonymous):

So 1/5 equals 1/10 ?

OpenStudy (tkhunny):

No. That is not correct. 1/5 = 2/10. How can you know that?

OpenStudy (anonymous):

I dont know, I was hoping somebody would explain it but also hint the answer at the same time.

OpenStudy (tkhunny):

You must know something about fractions. That's what we're doing, here - trying to fond out what you know about fractions. So, how many pieces did I eat?

OpenStudy (anonymous):

2

OpenStudy (tkhunny):

Good. How many pieces remain for someone else?

OpenStudy (anonymous):

8

OpenStudy (tkhunny):

Establish: (1/5) * 10 = ???

OpenStudy (anonymous):

1/5 * 10 = 2

OpenStudy (tkhunny):

One more, before we get back to Egypt. (What) * 10 = 8

OpenStudy (anonymous):

2/10

OpenStudy (tkhunny):

No, we already stated that (2/10) * 10 = (1/5) * 10 = 2 Try again.

OpenStudy (anonymous):

it would then be 8/10...sorry ;/

OpenStudy (tkhunny):

Perfect. (8/10)*10 = (4/5)*10 = 8 Okay, we have a land mass. 19/20 is desert. What ratio represents the amount of land that is habitable or cultivated?

OpenStudy (anonymous):

0.95 or 19:20

OpenStudy (tkhunny):

That's what 19/20 means, but that doesn't answer my question.

OpenStudy (anonymous):

Im confused with what you're asking.

OpenStudy (tkhunny):

19/20 represents the ratio or portion of Egypt that is desert. 0.95 or 19:20 What represents the ratio or portion of Egypt that is NOT desert?

OpenStudy (anonymous):

1/20 ???

OpenStudy (tkhunny):

Perfect. 1/20 = 0.05 = 1:20 Given: Desert = 19/20 of Egypt and Habitable = 1/20 of Egypt How can we calculate the total land mass in each classification?

OpenStudy (anonymous):

I thought we already did ? Sorry if im wasting time... @tkhunny

OpenStudy (tkhunny):

We absolutely did not do that. We cut up a pie. We have yet to cut up Egypt. How can we calculate the total land mass in each classification?

OpenStudy (anonymous):

Dividing ?

OpenStudy (tkhunny):

Maybe. How did we do the pie? 10*(1/5) = 2 10*(4/5) = 8 Does that look like division or multiplication?

OpenStudy (anonymous):

Multiplication..Thank you for being patient with someone like me.

OpenStudy (tkhunny):

I'm not complaining. I'm here to help. 995,450 mi2 * (WHAT) = Amount of Habitable Land 995,450 mi2 * (WHAT) = Amount of Desert

OpenStudy (anonymous):

995,450 mi2 * 1/20 = amount of habitable land 995, 450 mi2 * 19/20 = amount of desrt

OpenStudy (tkhunny):

Perfect. Can you now do the multiplication?

OpenStudy (tkhunny):

You should notice that 1 + 19 = 20. This is NOT a coincidence.

OpenStudy (anonymous):

49772.5 = 995,450 * 1/20 945677.5 = 955,450 * 19/20

OpenStudy (tkhunny):

Check your answer and see if it makes sense. 49772.5 + 945677.5 = ??? Do we get the original total back?

OpenStudy (anonymous):

No...

OpenStudy (tkhunny):

It's probably because you used 955, rather than 995 in one of the multiplications. Try again until the total comes back.

OpenStudy (anonymous):

i still get the same thing

OpenStudy (tkhunny):

Okay, let's try checking again. Do you get the original total? 995,450 * 1/20 = 49772.5 995,450 * 19/20 = 945677.5 49772.5 + 945677.5 =

OpenStudy (anonymous):

I get 995,450 so yes.

OpenStudy (tkhunny):

There we go, then. You should now have enough information to answer the questions from the beginning. Do them in a different order. Write the mathematical expression or expressions that you will use to solve this problem. Carry out the operations carefully. Show all of your work. Check your results. Explain, using complete sentences, how you will solve this problem.

OpenStudy (anonymous):

Thank you !! I really appreciate all your help <3 If I have anymore questions can I message you ?

OpenStudy (tkhunny):

Messaging is no good. Math expressions are too difficult in that environment. Just post to the public boards and tag me if you wish. I may still be here.

OpenStudy (anonymous):

Im stuck on Part D...how would I write it ? @tkhunny Solve: (5 points) Carry out the operations carefully. Show all of your work.

OpenStudy (tkhunny):

How did you answer the other parts?

OpenStudy (anonymous):

This is what I have so far; A) You are being asked to find the area of land that is desert and area of land that is usable. B) Total area of Egypt is 995,450 mi. 19/20 area is desert. I will solve this problem by multiplying and adding. C) 995,450 * 1/20 = 49772.5 995,450 * 19/20 = 945677.5 49772.5 = 945677.5 = 995,450)

OpenStudy (tkhunny):

Can't really help you with that. You need to describe your process. Give it a go. Make it make sense.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you sooooo much for all your help. I apologize if i was whiny or annoying or snappy

OpenStudy (tkhunny):

No worries. It's not supposed to be easy. :-)